# The Effect of Massive Trans-Neptunian Objects in the Long-term Evolution and Leaking Rates of Neptune’s 3:2 and 2:1 Mean Motion Resonances

### v2

@@ Abstract
software: This work has made use of the integrator package Rebound (H. Rein & S. F. Liu, 2012), and the Python modules Matplotlib (J. D. Hunter, 2007), and NumPy (C. R. Harris et al., 2020).
@@ 1 Introduction
Neptune’s mean motion resonances (MMRs) in the Kuiper belt (KB) are dynamical regions where a significant fraction of the total observed trans-Neptunian objects (TNOs) are located (e.g. P. S. Lykawka & T. Mukai, 2007; B. Gladman et al., 2012; M. Alexandersen et al., 2016; K. Volk et al., 2016). For example, in the full data release of the Origin of the Solar System Objects Survey (OSSOS, M. T. Bannister et al., 2018), 132 Plutinos are listed out of a total of 838 well-characterized TNOs [^1] From a total of 840 TNOs discovered in the OSSOS survey, only two objects were not tracked to high-precision orbits., while another 34 in the 2:1 (a.k.a. Twotinos, Y.-T. Chen et al., 2019), 39 objects in the 7:4, and 29 objects in the 5:2 MMRs (B. L. Crompvoets et al., 2022) were found in the survey; those numbers show that $\sim$27% of the total OSSOS-characterized discoveries were located in just four Neptune MMRs. More recently, E. Forgács-Dajka et al. (2023) performed a classification of Kuiper belt objects (KBOs) located in MMRs, finding that from the 4121 objects with semimajor axis $a>30.1$ au, listed in the JPL Horizons database [^2] https://ssd.jpl.nasa.gov/horizons/, 906 librate, for at least 100 Myr, in one of several possible MMRs with Neptune; their resonant objects were identified using their custom Fast Identification of MMRs method (or FAIR method, E. Forgács-Dajka et al., 2018). Based on these results, $\sim 22\%$ of the reported TNOs, observed by various surveys, are located within the MMRs with Neptune.
@@
Two of the most significant resonances associated with Neptune are the 3:2 and 2:1. This is important for several reasons. First, most well-characterized resonant objects belong to the 3:2 population, often referred to as “Plutinos”. This group includes Pluto, the most notable and well-studied trans-Neptunian object (TNO), both physically and dynamically (e.g. A. Milani et al., 1989; S. A. Stern et al., 2018). Additionally, the locations of the 3:2 and 2:1 resonances define the inner and outer boundaries of the main classical Kuiper Belt, respectively (see B. Gladman et al., 2008). Moreover, the orbital distribution of the 2:1 resonance population, known as “Twotinos”, could provide crucial evidence for understanding the characteristics of Neptune’s early migration (S. Ida et al., 2000; E. I. Chiang & A. B. Jordan, 2002; M. C. Wyatt, 2003; R. A. Murray-Clay & E. I. Chiang, 2005; Y.-T. Chen et al., 2019; H. Li & L.-Y. Zhou, 2023).
@@
Indeed, the populations of Neptune’s MMRs are thought to originate from the very early stages of solar system evolution, during the giant planets’ outward migration stage, due to the interaction of the giants with a massive disk of remaining planetesimals (J. A. Fernandez & W. H. Ip, 1984; D. Nesvorný, 2018). The first theory to explain the population of MMRs (and the origin of Pluto’s orbit in particular) is due to R. Malhotra (1993, 1995). In this framework, often referred to as ”adiabatic resonance sweeping”, the slow, smooth migration of Neptune carries its first-order MMRs through an external, dynamically cold planetesimal belt (including Pluto). Planetesimals captured into these resonances are transported outward along with Neptune, while their eccentricities—and, to a lesser extent, their inclinations—are adiabatically excited (e.g. R. Malhotra, 1998; R. S. Gomes, 2000).
@@
While this process sweeps an overly large number of objects into the MMRs, their eccentricities naturally increase, possibly bringing them close enough to Neptune or Uranus to eject them from the KB (e.g. A. Morbidelli, 1997; M. A. Muñoz-Gutiérrez et al., 2018). Recently, a non-smooth, but instead grainy migration of Neptune has been favoured to explain the relatively small number of particles trapped in the resonances (D. Nesvorný & D. Vokrouhlický, 2016; S. M. Lawler et al., 2019), as sudden jumps during the migration would release some of the trapped TNOs, making such a migration less efficient in trapping objects into the MMR.
@@
Once the planetary migration ends, a few to several dozens of Myr after the formation of the solar system, it is believed that the populations trapped in MMRs will evolve steadily, provided escape rates are low; this would imply, at the same time, that ratios between population sizes of different MMRs should only slowly evolve along the solar system’s age. (e.g. J. M. Hahn & R. Malhotra, 2005; M. S. Tiscareno & R. Malhotra, 2009).
@@
On the other hand, it is well known that a leaking process from MMRs has constantly occurred over the age of the solar system, either due to a weak chaotic diffusion present inside MMRs or by the direct perturbations of massive bodies inside and outside of the resonances (A. Morbidelli, 1997; W.-H. Ip & J. A. Fernandez, 1997; M. S. Tiscareno & R. Malhotra, 2009; M. A. Muñoz-Gutiérrez et al., 2019). In contrast to the above process, a refilling or resupplying of those same MMRs with new material, mainly coming from the classical KB, has not been studied nor considered important when estimating the long-term evolution of the populations trapped inside MMRs with Neptune. Nonetheless, the latter process could significantly affect the population ratios we observe today and obscure the optimistic constraints imposed by planetary migration models based on current resonant population numbers and distributions.
@@
In a previous work, we showed that dwarf-planetary-sized perturbers could contribute to the resupplying of MMRs in cold debris disks, increasing the injection rate of low-inclination comets (M. A. Muñoz-Gutiérrez et al., 2018). In the solar system, the 34 largest TNOs can increase the number of Jupiter Family Comets (JFCs) injected into the inner solar system (M. A. Muñoz-Gutiérrez et al., 2019). In that work, we did not observe significant variation in the evolution of the resonances, except for Pluto in the 3:2 MMR, where an increase in JFCs is evident (see Figure 5 in M. A. Muñoz-Gutiérrez et al., 2019). This indicates that more objects are lost from the 3:2 MMR when Pluto is present compared to when it is absent. Given that the then-known 34 most massive objects do not directly influence any other resonances, we question whether less massive objects, but closer to the resonances, might have a more substantial impact on their evolution. This work aims to establish limits on this question using high-resolution numerical simulations.
@@
This paper is organized as follows. Section 2 presents our two data sample sources. Section 3 describes our simulations and the initial characterization of potentially long-term resonant objects. Section 4 shows the behavior of the resonant populations, particularly their leaking rates as a function of time. Finally, our conclusions are presented in Section 5.
2 Data
@@
Name $H_{V}$ $a$ (au) p $\rho$ (g/cm3) RR(km) MM($\times 10^{-3}M_{\earth}$)
@@
a: Mass of a single object at the barycenter, considering Charon’s contribution. S. A. Stern et al. (2018)
b: Radius and density known from A. C. Barr & M. E. Schwamb (2016).
@@
c: Only radius known from E. Lellouch et al. (2013); density and mass derived from M. A. Muñoz-Gutiérrez et al. (2019).
d: Radius and density derived from M. Mommert et al. (2012); A. Dias-Oliveira et al. (2017).
@@
e: Only radius known from M. Mommert et al. (2012); density and mass derived from M. A. Muñoz-Gutiérrez et al. (2019).
f: All values derived using Eqs. 4 to 6 in M. A. Muñoz-Gutiérrez et al. (2019).
2.1 Initial Retrieval of Resonant Populations
@@
In this work, we explore the long-term evolution of the two strongest MMRs in the KB, namely the 3:2 and 2:1 first-order resonances with Neptune, which together account for approximately two-thirds of the observed population in all MMRs (B. Gladman et al., 2012; E. Forgács-Dajka et al., 2023). We used two different sources of data to define initial, potentially long-term, librating populations inside these resonances. The first set of data was obtained from observations and comes from NASA JPL’s Small-Body Data-Base (SBDB) [^3] https://ssd.jpl.nasa.gov/tools/sbdb_query.html, as well as from The International Astronomical Union Minor Planet Center (MPC) [^4] https://www.minorplanetcenter.net/data. For this observational sample, we retrieved data from NASA JPL’s SBDB, applying generous and straightforward criteria for the semimajor axis, $a$, selecting objects in the interval $38.8<a<40$ au for potential members of the 3:2 MMR and within $47<a<48.5$ au for potential members of the 2:1 MMR. These selection criteria yield the orbital parameters of 563 potential Plutinos (in the 3:2 MMR) and 153 potential Twotinos (in the 2:1 MMR).
@@
The second source of population data is the theoretical model obtained from the Canada - France Ecliptic Plane Survey (CFEPS), the so-called L7 model [^5] https://www.cfeps.net/?page_id=105< (J.-M. Petit et al., 2011; B. Gladman et al., 2012), which represents a debiased distribution of KBOs with magnitudes below $H_{g}\leq 8.5$; the L7 model lists objects in different families in the trans-Neptunian region, namely several MMRs, the classical belt, and the scattered disk.
@@
It is important to note that the L7 model is not intended to provide a comprehensive characterization of resonant populations. Instead, its goal is to reproduce the CFEPS observations after accounting for the well-understood survey biases, meaning it was designed to produce population estimates rather than serve as initial conditions for dynamical studies. However, given the L7 model’s clear ability to reasonably match observations, we have decided to consider this population as a potential source of resonant particles. This choice is supported by the understanding that the model provides a fairly acceptable representation of resonant populations. In any case, after a thorough initial characterization, we will only be left with particles that exhibit libration over an initial simulation period of 10 Myr, a procedure similar to what we perform in our previous work on the delivery of Jupiter-family comets (M. A. Muñoz-Gutiérrez et al., 2019).
@@
2.2 Orbital elements of the Giant Planets
@@
In this work, we performed a large number of numerical simulations with various initial test particle populations. Specifically, for the simulations using the L7 model to populate the resonances, we obtained precise heliocentric data for the giant planets (Jupiter, Saturn, Uranus, and Neptune) from JPL’s NASA Horizons system  [^6] https://ssd.jpl.nasa.gov/horizons/app.html, for the Julian Date (JD) JD2453157.5, which corresponds to June 1, 2004. Meanwhile, the SBDB objects are integrated starting at the epoch JD2460221, corresponding to October 3, 2023. For these simulations, we obtained precise heliocentric data for Jupiter, Saturn, Uranus, and Neptune from Horizons on the same date, JD2460221.
2.3 The Sample of the Ten Most Massive Plutinos and Twotinos
@@
From the observational sample, retrieved considering a large semimajor axis range, we identified the largest potentially resonant objects, which we will call prospective Plutinos and Twotinos, as those objects with the smallest absolute magnitudes. From these, we select the ten largest that, in an initial 10 Myr integration under the perturbation of the giant planets and the Sun, show consistent libration in either resonance.
Our list of the 20 largest objects can be found in Table 1. All these 20 identified objects coincided with the classification of K. Volk & C. Van Laerhoven (2024), i.e., they are classified as resonant based solely on 10 Myr integrations under perturbations from the Sun and the four GP. Among the set of the ten massive prospective Plutinos is, of course, Pluto, and other well-known Plutinos. Other objects in the prospective massive Plutino group, as well as all of the prospective massive Twotino group, are not so well characterized. This means that, from the 20 objects, only three of them have known masses: Pluto, Orcus, and $2003\,{\rm AZ}_{84}$. Of the other 17, two objects have known radius (Ixion and $2003\,{\rm VS}_{2}$), while for the other 15, only the absolute magnitude is known. To study the dynamical effect these objects would have on the MMR populations, we need to have at least an approximate determination of the mass. We do this using the fitting procedure of M. A. Muñoz-Gutiérrez et al. (2019), where unknown albedos are assigned as a function of absolute magnitude, $H_{V}$, as follows:
@@
Since all our objects without precise data have $H_{V}>3.21$, we assign a constant albedo of 0.140 to most of the objects in our list. We then obtain a radius following (A. W. Harris & A. W. Harris, 1997) as,
@@
Finally, we assign a density as a function of radius following again the fitting from M. A. Muñoz-Gutiérrez et al. (2019) given by:
@@
Equations 1 and 3 were determined by fitting the data of 27 objects with known albedo, and 13 objects with known density, respectively. For details about this procedure, see Appendix A in M. A. Muñoz-Gutiérrez et al. (2019).
These objects will be used as a set of massive perturbers in later simulations to investigate how these dwarf-planet-sized objects affect the leakage rates of the 3:2 and 2:1 MMRs.
Figure 1: Initial orbital parameters of retrieved objects in the 3:2 (left panels) and 2:1 (right panels) MMRs. The top panels show the distribution in semimajor axis vs eccentricity, while the lower panels show semimajor axis vs inclination. Gray dots show the sample obtained from the L7 model; blue dots show the observational sample obtained from JPL’s SBDB. An initial statistical characterization is shown in dashed lines, which represent the median values of semimajor axis (vertical dashed lines) and eccentricity and inclination (horizontal dashed lines) for the L7 and SBDB populations, color-coded in the same way as the dots.
2.4 Initial Orbital Distribution of Resonant Populations

### v3

@@ Abstract
software: This work has made use of the integrator package Rebound (Rein and Liu, 2012), and the Python modules Matplotlib (Hunter, 2007), and NumPy (Harris et al., 2020).
@@ I Introduction
Neptune’s mean motion resonances (MMRs) in the Kuiper belt (KB) are dynamical regions where a significant fraction of the total observed trans-Neptunian objects (TNOs) are located (e.g. Lykawka and Mukai, 2007; Gladman et al., 2012; Alexandersen et al., 2016; Volk et al., 2016). For example, in the full data release of the Origin of the Solar System Objects Survey (OSSOS; Bannister et al., 2018), 132 Plutinos are listed out of a total of 838 well-characterized TNOs [^1] From a total of 840 TNOs discovered in the OSSOS survey, only two objects were not tracked to high-precision orbits., while another 34 in the 2:1 (a.k.a. Twotinos; Chen et al., 2019), 39 objects in the 7:4, and 29 objects in the 5:2 MMRs (Crompvoets et al., 2022) were found in the survey; those numbers show that $\sim$27% of the total OSSOS-characterized discoveries were located in just four Neptune MMRs. More recently, Forgács-Dajka et al. (2023) performed a classification of Kuiper belt objects (KBOs) located in MMRs, finding that from the 4121 objects with semimajor axis $a>30.1$ au, listed in the JPL Horizons database [^2] https://ssd.jpl.nasa.gov/horizons/, 906 librate, for at least 100 Myr, in one of several possible MMRs with Neptune; their resonant objects were identified using their custom Fast Identification of MMRs method (or FAIR method, Forgács-Dajka et al., 2018). Based on these results, $\sim 22\%$ of the reported TNOs, observed by various surveys, are located within the MMRs with Neptune.
@@
Two of the most significant resonances associated with Neptune are the 3:2 and 2:1. This is important for several reasons. First, most well-characterized resonant objects belong to the 3:2 population, often referred to as “Plutinos”. This group includes Pluto, the most notable and well-studied trans-Neptunian object (TNO), both physically and dynamically (e.g. Milani et al., 1989; Stern et al., 2018). Additionally, the locations of the 3:2 and 2:1 resonances define the inner and outer boundaries of the main classical Kuiper Belt, respectively (see Gladman et al., 2008). Moreover, the orbital distribution of the 2:1 resonance population, known as “Twotinos”, could provide crucial evidence for understanding the characteristics of Neptune’s early migration (Ida et al., 2000; Chiang and Jordan, 2002; Wyatt, 2003; Murray-Clay and Chiang, 2005; Chen et al., 2019; Li and Zhou, 2023).
@@
Indeed, the populations of Neptune’s MMRs are thought to originate from the very early stages of solar system evolution, during the giant planets’ outward migration stage, due to the interaction of the giants with a massive disk of remaining planetesimals (Fernandez and Ip, 1984; Nesvorný, 2018). The first theory to explain the population of MMRs (and the origin of Pluto’s orbit in particular) is due to Malhotra (1993, 1995). In this framework, often referred to as “adiabatic resonance sweeping”, the slow, smooth migration of Neptune carries its first-order MMRs through an external, dynamically cold planetesimal belt (including Pluto). Planetesimals captured into these resonances are transported outward along with Neptune, while their eccentricities—and, to a lesser extent, their inclinations—are adiabatically excited (e.g. Malhotra, 1998; Gomes, 2000).
@@
While this process sweeps an overly large number of objects into the MMRs, their eccentricities naturally increase, possibly bringing them close enough to Neptune or Uranus to eject them from the KB (e.g. Morbidelli, 1997; Muñoz-Gutiérrez et al., 2018). Recently, a non-smooth, but instead grainy migration of Neptune has been favoured to explain the relatively small number of particles trapped in the resonances (Nesvorný and Vokrouhlický, 2016; Lawler et al., 2019), as sudden jumps during the migration would release some of the trapped TNOs, making such a migration less efficient in trapping objects into the MMR.
An additional possibility studied in the literature is the existence of a rogue planet present in the outer solar system during its early stages; such an object can enhance the capture efficiency of Neptune’s far MMRs (Huang et al., 2022). Although this scenario has only been studied farther out than the 2:1 MMRs, its efficiency could be of importance for closer resonances as well.
@@
Once the planetary migration ends, a few to several dozens of Myr after the formation of the Solar System, it is believed that the populations trapped in MMRs will evolve steadily, provided escape rates are low; this would imply, at the same time, that ratios between population sizes of different MMRs should only slowly evolve along the solar system’s age. (e.g. Hahn and Malhotra, 2005; Tiscareno and Malhotra, 2009).
@@
On the other hand, it is well known that a leaking process from MMRs has constantly occurred over the age of the Solar System, either due to a weak chaotic diffusion present inside MMRs or by the direct perturbations of massive bodies inside and outside of the resonances (Morbidelli, 1997; Ip and Fernandez, 1997; Tiscareno and Malhotra, 2009; Muñoz-Gutiérrez et al., 2019). In contrast to the above process, a refilling or resupplying of those same MMRs with new material, mainly coming from the classical KB, has not been studied nor considered important when estimating the long-term evolution of the populations trapped inside MMRs with Neptune. Nonetheless, the latter process could significantly affect the population ratios we observe today and obscure the optimistic constraints imposed by planetary migration models based on current resonant population numbers and distributions.
@@
In a previous work, we showed that dwarf-planetary-sized perturbers could contribute to the resupplying of MMRs in cold debris disks, increasing the injection rate of low-inclination comets (Muñoz-Gutiérrez et al., 2018). In the solar system, the 34 largest TNOs can increase the number of Jupiter Family Comets (JFCs) injected into the inner Solar System (Muñoz-Gutiérrez et al., 2019). In that work, we did not observe significant variation in the evolution of the resonances, except for Pluto in the 3:2 MMR, where an increase in JFCs is evident (see Figure 5 in Muñoz-Gutiérrez et al., 2019). This indicates that more objects are lost from the 3:2 MMR when Pluto is present compared to when it is absent. Given that the then-known 34 most massive objects do not directly influence any other resonances, we question whether less massive objects, but closer to the resonances, might have a more substantial impact on their evolution. This work aims to establish limits on this question using high-resolution numerical simulations.
@@
This paper is organized as follows. Section II presents our two data sample sources. Section III describes our simulations and the initial characterization of potentially long-term resonant objects. Section IV shows the behavior of the resonant populations, particularly their leaking rates as a function of time. Finally, our conclusions are presented in Section V.
II Data
@@
Name $H_{V}$ $a$ (au) p $\rho$ (g/cm3) RR(km) MM($\times 10^{-4}M_{\earth}$)
@@
a: Mass of a single object at the barycenter, considering Charon’s contribution. Stern et al. (2018)
b: Radius and density known from Barr and Schwamb (2016).
@@
c: Only radius known from Lellouch et al. (2013); density and mass derived from Muñoz-Gutiérrez et al. (2019).
d: Radius and density derived from Mommert et al. (2012); Dias-Oliveira et al. (2017).
@@
e: Only radius known from Mommert et al. (2012); density and mass derived from Muñoz-Gutiérrez et al. (2019).
f: All values derived using Eqs. 4 to 6 in Muñoz-Gutiérrez et al. (2019).
II.1 Initial Retrieval of Resonant Populations
@@
In this work, we explore the long-term evolution of the two strongest MMRs in the KB, namely the 3:2 and 2:1 first-order resonances with Neptune, which together account for approximately two-thirds of the observed population in all MMRs (Gladman et al., 2012; Forgács-Dajka et al., 2023). We used two different sources of data to define initial, potentially long-term, librating populations inside these resonances. The first set of data was obtained from observations and comes from NASA JPL’s Small-Body Data-Base (SBDB) [^3] https://ssd.jpl.nasa.gov/tools/sbdb_query.html, as well as from The International Astronomical Union Minor Planet Center (MPC) [^4] https://www.minorplanetcenter.net/data. For this observational sample, we retrieved data from NASA JPL’s SBDB, applying generous and straightforward criteria for the semimajor axis, $a$, selecting objects in the interval $38.8<a<40$ au for potential members of the 3:2 MMR and within $47<a<48.5$ au for potential members of the 2:1 MMR. These selection criteria yield the orbital parameters of 563 potential Plutinos (in the 3:2 MMR) and 153 potential Twotinos (in the 2:1 MMR).
@@
The second source of population data is the theoretical model obtained from the Canada - France Ecliptic Plane Survey (CFEPS), the so-called L7 model [^5] https://www.cfeps.net/?page_id=105< (Petit et al., 2011; Gladman et al., 2012), which represents a debiased distribution of KBOs with magnitudes below $H_{g}\leq 8.5$; the L7 model lists objects in different families in the trans-Neptunian region, namely several MMRs, the classical belt, and the scattered disk.
@@
It is important to note that the L7 model is not intended to provide a comprehensive characterization of resonant populations. Instead, its goal is to reproduce the CFEPS observations after accounting for the well-understood survey biases, meaning it was designed to produce population estimates rather than serve as initial conditions for dynamical studies. However, given the L7 model’s clear ability to reasonably match observations, we have decided to consider this population as a potential source of resonant particles. This choice is supported by the understanding that the model provides a fairly acceptable representation of resonant populations. In any case, after a thorough initial characterization, we will only be left with particles that exhibit libration over an initial simulation period of 10 Myr, a procedure similar to what we perform in our previous work on the delivery of Jupiter-family comets (Muñoz-Gutiérrez et al., 2019).
@@
II.2 Orbital elements of the Giant Planets
@@
In this work, we performed a large number of numerical simulations with various initial test particle populations. Specifically, for the simulations using the L7 model to populate the resonances, we obtained precise heliocentric data for the giant planets (Jupiter, Saturn, Uranus, and Neptune) from JPL’s NASA Horizons system  [^6] https://ssd.jpl.nasa.gov/horizons/app.html, for the Julian Date (JD) JD2453157.5, which corresponds to June 1, 2004, the epoch used for the L7 model generation, which lies at the middle of the CFEPS survey. Meanwhile, the SBDB objects are integrated starting at the epoch JD2460221, corresponding to October 3, 2023. For these simulations, we obtained precise heliocentric data for Jupiter, Saturn, Uranus, and Neptune from Horizons on the same date, JD2460221.
II.3 The Sample of the Ten Most Massive Plutinos and Twotinos
@@
From the observational sample, retrieved considering a large semimajor axis range, we identified the largest potentially resonant objects, which we will call prospective Plutinos and Twotinos, as those objects with the smallest absolute magnitudes. From these, we select the ten smallest $H_{V}$ objects that, in an initial 10 Myr integration under the perturbation of the giant planets and the Sun, show consistent libration in either resonance.
Our list of the 20 largest objects can be found in Table 1. All these 20 identified objects coincided with the classification of Volk and Van Laerhoven (2024), i.e., they are classified as resonant based solely on 10 Myr integrations under perturbations from the Sun and the four GP. Among the set of the ten massive prospective Plutinos is, of course, Pluto, and other well-known Plutinos. Other objects in the prospective massive Plutino group, as well as all of the prospective massive Twotino group, are not so well characterized. This means that, from the 20 objects, only three of them have known masses: Pluto, Orcus, and $2003\,{\rm AZ}_{84}$. Of the other 17, two objects have known radius (Ixion and $2003\,{\rm VS}_{2}$), while for the other 15, only the absolute magnitude is known. To study the dynamical effect these objects would have on the MMR populations, we need to have at least an approximate determination of the mass. We do this using the fitting procedure of Muñoz-Gutiérrez et al. (2019), where unknown albedos are assigned as a function of absolute magnitude, $H_{V}$, as follows:
@@
Since all our objects without precise data have $H_{V}>3.21$, we assign a constant albedo of 0.140 to most of the objects in our list. We then obtain a radius following (Harris and Harris, 1997) as,
@@
Finally, we assign a density as a function of radius following again the fitting from Muñoz-Gutiérrez et al. (2019) given by:
@@
Equations 1 and 3 were determined by fitting the data of 27 objects with known albedo, and 13 objects with known density, respectively. For details about this procedure, see Appendix A in Muñoz-Gutiérrez et al. (2019).
These objects will be used as a set of massive perturbers in later simulations to investigate how the largest resonant objects affect the leakage rates of the 3:2 and 2:1 MMRs.
Figure 1: Initial orbital parameters of retrieved objects in the 3:2 (left panels) and 2:1 (right panels) MMRs. The top panels show the distribution in semimajor axis vs. eccentricity, while the lower panels show semimajor axis vs inclination. Gray dots show the sample obtained from the L7 model; blue dots show the observational sample obtained from JPL’s SBDB. An initial statistical characterization is shown in dashed lines, which represent the median values of semimajor axis (vertical dashed lines) and eccentricity and inclination (horizontal dashed lines) for the L7 and SBDB populations, color-coded in the same way as the dots.
II.4 Initial Orbital Distribution of Resonant Populations